OpenStudy (anonymous):

Please help me to solve this mathematical induction problem, 2^2+5^2+8^2+….+(3n-1)^2=1n/2(6n^2+3n-1)

OpenStudy (anonymous):

Have you gone through the steps (e.g. assume true for k...)

OpenStudy (anonymous):

Proof: For each positive integer \(n\ge1\) , let \(S(n)\) be the statement \[2^2+5^2+8^2+...+(3n-1)^2=\frac{n}{2}(6n^2+3n-1)\] Basis step: \(S(1)\) is the statement \(2^2=\frac{1}{2}(6+3-1)=4\). Thus \(S(1)\) is true. Inductive step: We suppose that \(S(k)\) is true and prove that \(S(k+1)\) is true. Thus, we assume that \[2^2+5^2+8^2+...+(3k-1)^2=\frac{k}{2}(6k^2+3k-1)\] and prove that \[2^2+5^2+8^2+...+(3k+1)^2=\frac{k+1}{2}(6(k+1)^2+3(k+1)-1)\]

OpenStudy (zzr0ck3r):

lol, I dont see how any of this helps

OpenStudy (zzr0ck3r):

but it looks cool

OpenStudy (anonymous):

me too i dont see...lol :)

OpenStudy (anonymous):

ok @nroprstar u need to show that\[\frac{k}{2}(6k^2+3k-1)+(3k+1)^2=\frac{k+1}{2}(6(k+1)^2+3(k+1)-1)\]why and how??? this is ur work !! :)

OpenStudy (anonymous):

oops thats\[\frac{k}{2}(6k^2+3k-1)+(3k+2)^2=\frac{k+1}{2}(6(k+1)^2+3(k+1)-1)\]